what is the width of rectangle with a perimeter of 70 ft if its lenghth is 1 foot
Answer by checkley77(12844)

2L+2W=70
2*1+2W=70
2+2W=70
2W=70-2
2W=68
W=68/2
W=34 ANS. FOR THE WIDTH.
PROOF:
2*1+2*34=70
2+68=70
70=70